OpenStudy (anonymous):

In degrees, what are all values of 'x' between 0 and 360 for which \[\sin x > \sqrt{1-\sin^2 x}\]?

OpenStudy (anonymous):

right hand side is \[|\cos(x)|\] if that helps

OpenStudy (anonymous):

i believe your answer (in degrees) is 45 degrees to 135 decrees

OpenStudy (anonymous):

don't forget that \[\sqrt{1-\sin^2(x)}=|\cos(x)|\] and not \[\cos(x)\]

OpenStudy (anonymous):

how does that lead you to the solution?

OpenStudy (anonymous):

satellite is correct in using the absolute value....

OpenStudy (anonymous):

how do you square both sides in inequality?

OpenStudy (anonymous):